# A try for dark energy in quantum field theory: The vacuum energy of neutrino field

### v4

@@ IV Conclusion and Discussion
In this paper, the vacuum energy’s power-law divergence has been established. The physical result in the UV-free scheme indicates that it is not a problem in the UV region. Freed from the UV divergence covering, the actual question is that contributions of heavy fields should be suppressed by unknown mechanisms. Here we give more interpretations about two common questions as mentioned by Ref. Cline:2024klo , i.e., why we choose the value of $\mu_{\Lambda}$ in meV scale, and why we ignore the heavy fields that could contribute. To the first question about the value of $\mu_{\Lambda}$, it is currently lacking pathways to derive its specific value from fundamental principles (e.g., the values of particle masses and coupling constants in SM). Here the scale $\mu_{\Lambda}$ is taken as an input parameter, and its parameter space is derived by the observation of dark energy density with naturalness. This energy scale preserves the spacetime flatness of the present-day universe. Theoretically, this scale may link to the inflationary field or neutrino mass origin, etc. To the second question, this work explores the $\mu_{\Lambda}-$characterized scale decoupling of vacuum energy from microscopic quantity to macroscopic effect, implying heavy fields remain inert. The answer to it is a long exploring journey.

### v5

@@ IV Conclusion and Discussion
In this paper, the power-law divergence of vacuum energy is established. The physical result obtained in the UV-free scheme indicates it is not a problem in the UV region. Freed from the veil of UV divergence, the core question becomes how contributions of heavy fields are suppressed by unknown mechanisms, constituting a new open question. This work thereby recasts the cosmological constant problem from a fine-tuning puzzle into a defined physical question, thus clarifying the actual nature of the problem. In fundamental physics, such a redefinition is itself an advance, while its full resolution is the next open goal. A new framework comes with its necessary transformation of perspective. Accordingly, we address two recurrent questions (also noted in Ref. Cline:2024klo ): why the value of $\mu_{\Lambda}$ is chosen at the meV scale, and why contributions from heavy fields are neglected. To the first question about the value of $\mu_{\Lambda}$, it is currently lacking pathways to derive its specific value from fundamental principles (e.g., the values of particle masses and coupling constants in SM). Here $\mu_{\Lambda}$ is taken as an input parameter, and its parameter space is derived by the observation of dark energy density with naturalness. Theoretically, this scale may link to the inflationary field or neutrino mass origin, etc. A connection to neutrinos is of particular interest, as their mass scale would interface quantum vacuum physics with the cosmos. To the second question, this work explores the $\mu_{\Lambda}$-characterized scale decoupling of vacuum energy from microscopic quantity to macroscopic effect, implying heavy fields remain inert. This suggests that the dark energy problem could be minimally addressed by introducing the new scale $\mu_{\Lambda}$ into SM.
@@ Acknowledgments
Appendix A Fluctuation factor
For a mass $m$ carrier field, the $\mu_{\Lambda}$-characterized fluctuation factor describes macroscopic effects from cumulative vacuum fluctuations. Its potential form is derived as follows. The first is basic assumptions, i.e.,
(a) Independence assumption. Consider a system with $N$ independent random fluctuations $\{q_{i}\}_{i=1}^{N}$, satisfying: Zero expectation, with $E[q_{i}]=0$; Finite variance, with $Var(q_{i})=\sigma_{i}^{2}$ (bounded fluctuation strength); Statistical independence between random variables $q_{i}$.
(b) Linear superposition. The total cumulative effect $X$ is a linear superposition of individual fluctuations with $X=\sum_{i=1}^{N}a_{i}q_{i}$, where $a_{i}$ being weight coefficients, and the system exhibits no significant nonlinear coupling.
Next is the derivation via the characteristic function. For each $q_{i}$, its characteristic function $\phi_{i}(\tau)$ = $E[e^{i\tau q_{i}}]$ expands to $\phi_{i}(\tau)$ = $1+i\tau E[q_{i}]-\frac{\tau^{2}}{2}E[q_{i}^{2}]+\cdots\approx 1-\frac{\sigma_{i}^{2}\tau^{2}}{2}$, where $\cdots$ denotes higher-order terms. By independence, the characteristic function of $X$ is the product of individual characteristic functions, $\phi_{X}(\tau)=\prod_{i=1}^{N}\phi_{i}(a_{i}\tau)\approx\prod_{i=1}^{N}(1-\frac{a_{i}^{2}\sigma_{i}^{2}\tau^{2}}{2})$. Taking the logarithm and linearizing (using $\ln(1+x)\approx x$ for $x\to 0$), the result is $\ln\phi_{X}(\tau)\approx-\frac{\tau^{2}}{2}\sum_{i=1}^{N}a_{i}^{2}\sigma_{i}^{2}$ $\Rightarrow$ $\phi_{X}(\tau)=\exp(-\frac{\sigma_{X}^{2}\tau^{2}}{2})$, where $\sigma_{X}^{2}=\sum_{i=1}^{N}a_{i}^{2}\sigma_{i}^{2}$. This is the characteristic function of a Gaussian distribution. In this paper, we assume that the cumulative effect $X$ of vacuum fluctuations manifests macroscopic effects through carrier fields. Specifically, the fluctuation factor $f(m)$ is set by $\phi_{X}(\tau)$, and it can be obtained via inverse Fourier transform of $\phi_{X}(\tau)$ with a matching condition $X\to m$, that is, $f(m)=A\frac{1}{2\pi}\int_{-\infty}^{\infty}e^{-i\tau m}\phi_{X}(\tau)\,d\tau=\frac{A}{\sqrt{2\pi\sigma_{X}^{2}}}\exp(-\frac{m^{2}}{2\sigma_{X}^{2}})=B\exp(-\frac{m^{2}\pi B^{2}}{A^{2}})$, with $B=\frac{A}{\sqrt{2\pi\sigma_{X}^{2}}}$. Considering photon field as the carrier field for normalization, one has $f(0)=1$, and thus $B=1$ is obtained. Taking $A^{2}/\pi=2\sigma_{X}^{2}=\frac{1}{k}\mu_{\Lambda}^{2}$, the fluctuation factor is written as $f(m)=e^{-km^{2}/\mu_{\Lambda}^{2}}$, and $k\sim 1$ can be adopted for $\mu_{\Lambda}$-characterized fluctuations.
From another perspective, the macroscopic effect of physical vacuum energy requires a long-range description with characteristic scale $\sim 1/\mu_{\Lambda}$. This reflects the filtering of microscopic energy into a macroscopic effective quantity, a principle common to statistical physics. A classic example is gas pressure, which stems not from each molecule’s rest energy $mc^{2}$, but from the small fraction in statistically averaged translational kinetic energy. Thus, macroscopic observables are only scale-selected subsets of microscopic energies.
In this analogy, the dark energy density is a macroscopic observable governed by a $\mu_{\Lambda}$-scale filter. The microscopic zero-point energy of a quantum field is enormous, yet only fluctuation modes coherent over $\sim 1/\mu_{\Lambda}$ contribute effectively. For a field with $m\gg\mu_{\Lambda}$, its coherence length (the Compton wavelength $\lambda_{c}\!\sim\!1/m$) is far smaller than $1/\mu_{\Lambda}$. Consequently, a macroscopic volume of size $\sim 1/\mu_{\Lambda}$ contains many independent fluctuation patches with random phases; their contributions cancel under macroscopic averaging via a random-walk process, leading to exponential suppression. Thus, the cumulative effect of a heavy field is strongly suppressed. The observed small dark energy density therefore does not contradict huge microscopic energies, but shows the cosmos is only sensitive to modes passing the decoupling filter $\lambda_{c}\gtrsim 1/\mu_{\Lambda}$. This phase-statistical filtering provides a statistical explanation for the smallness of dark energy density.